# Nonlinear transport theory at the order of quantum metric

### v1

@@ Abstract
Introduction.– The nonlinear Hall effect is characterized by a transverse voltage nonlinearly dependent on the longitudinal driving current. It has received tremendous attention because of its potential as a measurement method and for device applications Sodemann and Fu (2015); Low et al. (2015); Ma et al. (2019); Du et al. (2018); Kang et al. (2019); Du et al. (2019); Nandy and Sodemann (2019); Xiao et al. (2019); Isobe et al. (2020); Du et al. (2021a, b); Kumar et al. (2021); Ho et al. (2021); He et al. (2021); Watanabe and Yanase (2020); Tiwari et al. (2021); König and Levchenko (2021); Duan et al. (2022); He et al. (2022); Huang et al. (2022); Kaplan et al. (2023); Min et al. (2023); Kaplan et al. (2024); Chen et al. (2024a); Wang et al. (2024); Xiang and Wang (2024); Chen et al. (2024b); Yao et al. (2024); He et al. (2024). More excitingly, without time-reversal and inversion symmetries but under a combined time-reversal and inversion symmetry, the nonlinear Hall effect can be contributed by the quantum metric Gao et al. (2014); Wang et al. (2021); Liu et al. (2021); Gao et al. (2023); Wang et al. (2023); Fang et al. (2024), the real part of quantum geometry Provost and Vallee (1980); Resta (2011); Törmä et al. (2022); Törmä (2023); Bouhon et al. (2023), which measures the distance between quantum states, thus can provide spacetime insight to the challenges in condensed matter Liu et al. (2024). The nonlinear Hall effect from the quantum metric dipole ($\chi^{\mathrm{QMD}}$ and $\mathcal{G}_{l}^{ab}$ in Tab. 1) is essentially an anomalous Hall effect but with the driving current acting twice. In this sense, the disorder contributions, such as the side-jump and skew-scattering Nagaosa et al. (2010), cannot be ignored Du et al. (2019), inspiring an increasing interest in exhausting all contributions to the nonlinear transport Ma et al. (2023); Atencia et al. (2023).
@@
Table 1: Formulas of the quadratic nonlinear conductivity at the order of quantum metric, including the quantum metric dipole (QMD), Drude, mixed Berry curvature and skew-scattering (BSK), mixed Berry curvature and side-jump (BSJ), second-order side-jump (2SJ) and skew-scattering (2SK), mixed side-jump and skew-scattering (SSK). $\tau$ is the scattering time, $-e$ is the electron charge, $\varepsilon^{abc}$ is the anti-symmetric tensor, Fermi distribution $f_{l}^{(0)}=1/[\exp[(\epsilon_{l}-\epsilon_{F})/k_{B}T]+1]$ with the Fermi energy $\epsilon_{F}$, $l$ stands for the band index $\gamma$ and wave vector $\mathbf{k}=|k|e^{i\phi}$, $\Omega^{a}_{l}$ is the Berry curvature Xiao et al. (2010), $\mathcal{G}^{ab}_{l}$ is the Berry connection polarizability tensor, which is related to the quantum metric tensor $\mathrm{Re}(A^{\gamma\gamma^{\prime}}_{a}A^{\gamma^{\prime}\gamma}_{b})$ for two-band cases Gao et al. (2014); Wang et al. (2021); Liu et al. (2021), Berry connection $A^{\gamma\gamma^{\prime}}_{a}\equiv i\langle\gamma|\partial_{\mathbf{k}}^{a}|\gamma^{\prime}\rangle$ Xiao et al. (2010), $\partial^{b}_{\mathbf{k}}$$\equiv$$\partial/\partial k_{b}$, $\hat{\mathcal{H}}$ is the Hamiltonian, $\epsilon_{l}$ is the eigen energy, group velocity $v_{l}^{a}$ = $(1/\hbar)\partial\epsilon_{l}/\partial k_{a}$, $\partial_{\epsilon_{l}}\equiv\partial/\partial\epsilon_{l}$, side-jump velocity $v_{l}^{sj,a}=\sum_{l^{\prime}}\varpi^{S}_{ll^{\prime}}\delta r^{a}_{l^{\prime}l}$, $\varpi^{S}_{ll^{\prime}}$ ($\varpi_{l^{\prime}l}^{A}$) is the symmetric (anti-symmetric) scattering rate between states $l$ and $l^{\prime}$ (Sec. SIB of Sup ), $\delta r^{a}_{ll^{\prime}}$ is the coordinates shift with $V_{ll^{\prime}}$=$\langle l|\hat{V}_{imp}|l^{\prime}\rangle$ Du et al. (2019); Sinitsyn (2008), where the $\delta$-correlated random potential $\hat{V}_{imp}(\mathbf{r})$=$\sum_{i}V_{i}\delta(\mathbf{r}-\mathbf{R}_{i})$ can give the Gaussian $\langle V^{2}_{i}\rangle_{dis}$=$V^{2}_{0}$ and non-Gaussian correlations $\langle V^{3}_{i}\rangle_{dis}$=$V^{3}_{1}$ Du et al. (2019, 2021b); Sinitsyn (2008), leading to the Gaussian (intrinsic skew-scattering) $\varpi_{l^{\prime}l}^{(4A)}$ and non-Gaussian (extrinsic skew-scattering) $\varpi_{l^{\prime}l}^{(3A)}$ parts to $\varpi_{l^{\prime}l}^{A}$. $\bar{O}_{ll^{\prime}}^{a}$ ($\tilde{O}_{ll^{\prime}}^{a}$) is the symmetric (anti-symmetric) part of $O^{a}_{ll^{\prime}}$ Du et al. (2019), and $T_{ll^{\prime}}$ is the T-matrix Mahan (1990) (details in Secs. SI and SII of Sup ).
@@
$\mathcal{G}^{ab}_{l}\equiv 2e\mathrm{Re}\sum_{\gamma^{\prime}\neq\gamma}A^{\gamma\gamma^{\prime}}_{a}A^{\gamma^{\prime}\gamma}_{b}/(\varepsilon^{\gamma}_{\mathbf{k}}-\varepsilon^{\gamma^{\prime}}_{\mathbf{k}})$, $\chi_{1}^{\mathrm{SSK}}=\frac{\tau^{2}e^{3}}{2\hbar}\sum_{ll^{\prime}}\left(v_{l}^{a}+v_{l^{\prime}}^{a}\right)\tilde{O}_{l^{\prime}l}^{c}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$,
$\Omega^{a}_{l}\equiv$ $-2\varepsilon^{abc}\sum_{\gamma^{\prime}\neq\gamma}$ $\mathrm{Im}\langle\gamma|\partial^{b}_{\mathbf{k}}\hat{\mathcal{H}}|\gamma^{\prime}\rangle$ $\langle\gamma^{\prime}|\partial^{c}_{\mathbf{k}}\hat{\mathcal{H}}|\gamma\rangle$/$(\varepsilon^{\gamma}_{\mathbf{k}}-\varepsilon^{\gamma^{\prime}}_{\mathbf{k}})^{2}$, $\chi_{2}^{\mathrm{SSK}}=\frac{\tau^{3}e^{3}}{2\hbar}\sum_{ll^{\prime}l^{\prime\prime}}\left(v_{l^{\prime\prime}}^{a}+v_{l^{\prime}}^{a}\right)\varpi_{ll^{\prime}}^{A}\bar{O}_{l^{\prime\prime}l^{\prime}}^{c}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$,
@@
$O^{a}_{ll^{\prime}}$$\equiv$$(2\pi/\hbar)|T_{ll^{\prime}}|^{2}\delta r^{a}_{ll^{\prime}}\partial_{\epsilon_{l}}\delta(\epsilon_{l}-\epsilon_{l^{\prime}})$, $\quad\quad\quad+\frac{\tau^{3}e^{3}}{2\hbar}\sum_{ll^{\prime}}v_{l^{\prime}}^{a}v_{l}^{sj,b}\left(\partial_{\mathbf{k}}^{c}-\partial_{\mathbf{k}^{\prime}}^{c}\right)\varpi_{l^{\prime}l}^{A}\partial_{\varepsilon_{l}}f_{l}^{(0)}$,
$1/\tau\approx 1/\tau_{\bf k}\equiv\sum_{\bf k^{\prime}}\varpi_{\mathbf{k}^{\prime}\mathbf{k}^{\prime}}^{(2)}\left[1-\cos\left(\phi-\phi^{\prime}\right)\right]\delta\left(\varepsilon_{\mathrm{F}}-\epsilon_{l}\right)$. $\chi_{4}^{\mathrm{SSK}}=\frac{\tau^{3}e^{3}}{2\hbar^{2}}\sum_{ll^{\prime}}v_{l^{\prime}}^{sj,a}\left(\partial_{\mathbf{k}}^{c}-\partial_{\mathbf{k}^{\prime}}^{c}\right)\varpi_{l^{\prime}l}^{A}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$.
@@
Table 2: Nonzero quadratic conductivity elements $\chi_{abb}$ on the $x$-$y$ plane (defined as $J_{a}$ = $\chi_{abb}E_{b}^{2}$, $\{a,b\}\in\{x,y\}$, $abb$ is short for $\chi_{abb}$) at the order of quantum metric (all formulas in Tab. 1), can be found for only 9 out of the 122 magnetic point groups. The quantum metric dipole (QMD) Gao et al. (2014); Wang et al. (2021); Liu et al. (2021) as well as BSJ and BSK in this work have the anti-symmetric constraint $\chi_{abc}$ = $-\chi_{cba}$ Du et al. (2021b), so they have less nonzero elements than 2SJ, 2SK, and SSK found in this work Zhang et al. (2023). The $C_{2}$ axis is assumed in order as the $y$- and $x$- axis, the mirror plane $\sigma_{v}$ is assumed in order as the $zx$- and $yz$-plane, and the mirror plane $\sigma_{h}$ is assumed as the $x$-$y$ plane. We use the MSGCorep codes Liu et al. (2023b) to filter the groups without time-reversal and inversion but with the combined time-reversal and inversion symmetry and find their matrices.
@@
is large ($n_{i}$ is small, i.e., the clean limit), where $n_{i}$ is the impurity density, $V_{0}^{2}$ ($V_{1}^{3}$) measures the correlations of two (three) scattering events. $\chi^{\mathrm{QMD}}$, $\chi^{\mathrm{BSJ}}$, $\chi^{\mathrm{BSK}}_{2}$, and $\chi^{\mathrm{2SJ}}_{1,2,3}$ do not depend on $\lambda$, $\chi^{\mathrm{BSK}}_{1}$ and $\chi^{\mathrm{SSK}}_{2,3,4}$ are linear in $\lambda$, $\chi^{\mathrm{SSK}}_{2,3,4}$ has terms $\propto\lambda^{0}$ and $\lambda^{1}$, $\chi^{\mathrm{2SK}}_{1,2}$ has terms $\propto\lambda^{0}$, $\lambda^{1}$, and $\lambda^{2}$ (Tab. S5 of Sup ). The Drude contribution Sodemann and Fu (2015); Wang et al. (2021) depends on $1/(n_{i}V_{0}^{2})^{2}$. Also, for $\chi_{yxy}$ of a specific model, the disorder-induced but disorder-strength-independent contributions could overwhelms the quantum metric part Atencia et al. (2023).

### v2

@@ Abstract
Introduction.– The nonlinear Hall effect is characterized by a transverse voltage nonlinearly dependent on the longitudinal driving current. It has received tremendous attention because of its potential as a measurement method and for device applications Sodemann and Fu (2015); Low et al. (2015); Ma et al. (2019); Du et al. (2018); Kang et al. (2019); Du et al. (2019); Nandy and Sodemann (2019); Xiao et al. (2019); Isobe et al. (2020); Du et al. (2021a, b); Kumar et al. (2021); Ho et al. (2021); He et al. (2021); Watanabe and Yanase (2020); Tiwari et al. (2021); König and Levchenko (2021); Duan et al. (2022); He et al. (2022); Huang et al. (2022); Kaplan et al. (2023); Min et al. (2023); Kaplan et al. (2024); Chen et al. (2024a); Wang et al. (2024); Xiang and Wang (2024); Chen et al. (2024b); Yao et al. (2024); He et al. (2024). More excitingly, without time-reversal and inversion symmetries but under a combined time-reversal and inversion symmetry, the nonlinear Hall effect can be contributed by the quantum metric Gao et al. (2014); Wang et al. (2021); Liu et al. (2021); Gao et al. (2023); Wang et al. (2023); Fang et al. (2024), the real part of quantum geometry Provost and Vallee (1980); Resta (2011); Törmä et al. (2022); Törmä (2023); Hetényi and Lévay (2023); Bouhon et al. (2023), which measures the distance between quantum states, thus can provide spacetime insight to the challenges in condensed matter Liu et al. (2024). The nonlinear Hall effect from the quantum metric dipole ($\chi^{\mathrm{QMD}}$ and $\mathcal{G}_{l}^{ab}$ in Tab. 1) is essentially an anomalous Hall effect but with the driving current acting twice. In this sense, the disorder contributions, such as the side-jump and skew-scattering Nagaosa et al. (2010), cannot be ignored Du et al. (2019), inspiring an increasing interest in exhausting all contributions to the nonlinear transport Ma et al. (2023); Atencia et al. (2023).
@@
Table 1: Formulas of the quadratic nonlinear conductivity at the order of quantum metric, including the quantum metric dipole (QMD), Drude, mixed Berry curvature and skew-scattering (BSK), mixed Berry curvature and side-jump (BSJ), second-order side-jump (2SJ) and skew-scattering (2SK), mixed side-jump and skew-scattering (SSK). $\tau$ is the scattering time, $-e$ is the electron charge, $\varepsilon^{abc}$ is the anti-symmetric tensor, Fermi distribution $f_{l}^{(0)}=1/[\exp[(\epsilon_{l}-\epsilon_{F})/k_{B}T]+1]$ with the Fermi energy $\epsilon_{F}$, $l$ stands for the band index $\gamma$ and wave vector $\mathbf{k}=|k|\left(\cos\phi,\sin\phi\right)$ in 2D, $\Omega^{a}_{l}$ is the Berry curvature Xiao et al. (2010), $\mathcal{G}^{ab}_{l}$ is the Berry connection polarizability tensor, which is related to the quantum metric tensor $\mathrm{Re}(A^{\gamma\gamma^{\prime}}_{a}A^{\gamma^{\prime}\gamma}_{b})$ for two-band cases Gao et al. (2014); Wang et al. (2021); Liu et al. (2021), Berry connection $A^{\gamma\gamma^{\prime}}_{a}\equiv i\langle\gamma|\partial_{\mathbf{k}}^{a}|\gamma^{\prime}\rangle$ Xiao et al. (2010), $\partial^{b}_{\mathbf{k}}$$\equiv$$\partial/\partial k_{b}$, $\hat{\mathcal{H}}$ is the Hamiltonian, $\epsilon_{l}$ is the eigen energy, group velocity $v_{l}^{a}$ = $(1/\hbar)\partial\epsilon_{l}/\partial k_{a}$, $\partial_{\epsilon_{l}}\equiv\partial/\partial\epsilon_{l}$, side-jump velocity $v_{l}^{sj,a}=\sum_{l^{\prime}}\varpi^{S}_{ll^{\prime}}\delta r^{a}_{l^{\prime}l}$, $\varpi^{S}_{ll^{\prime}}$ ($\varpi_{l^{\prime}l}^{A}$) is the symmetric (anti-symmetric) scattering rate between states $l$ and $l^{\prime}$ (Sec. SIB of Sup ), $\delta r^{a}_{ll^{\prime}}$ is the coordinates shift with $V_{ll^{\prime}}$=$\langle l|\hat{V}_{imp}|l^{\prime}\rangle$ Du et al. (2019); Sinitsyn (2008), where the $\delta$-correlated random potential $\hat{V}_{imp}(\mathbf{r})$=$\sum_{i}V_{i}\delta(\mathbf{r}-\mathbf{R}_{i})$ can give the Gaussian $\langle V^{2}_{i}\rangle_{dis}$=$V^{2}_{0}$ and non-Gaussian correlations $\langle V^{3}_{i}\rangle_{dis}$=$V^{3}_{1}$ Du et al. (2019, 2021b); Sinitsyn (2008), leading to the Gaussian (intrinsic skew-scattering) $\varpi_{l^{\prime}l}^{(4A)}$ and non-Gaussian (extrinsic skew-scattering) $\varpi_{l^{\prime}l}^{(3A)}$ parts to $\varpi_{l^{\prime}l}^{A}$. $\bar{O}_{ll^{\prime}}^{a}$ ($\tilde{O}_{ll^{\prime}}^{a}$) is the symmetric (anti-symmetric) part of $O^{a}_{ll^{\prime}}$ Du et al. (2019), and $T_{ll^{\prime}}$ is the T-matrix Mahan (1990) (details in Secs. SI and SII of Sup ).
@@
$\mathcal{G}^{ab}_{l}\equiv 2e\mathrm{Re}\sum_{\gamma^{\prime}\neq\gamma}A^{\gamma\gamma^{\prime}}_{a}A^{\gamma^{\prime}\gamma}_{b}/(\epsilon^{\gamma}_{\mathbf{k}}-\epsilon^{\gamma^{\prime}}_{\mathbf{k}})$, $\chi_{1}^{\mathrm{SSK}}=\frac{\tau^{2}e^{3}}{2\hbar}\sum_{ll^{\prime}}\left(v_{l}^{a}+v_{l^{\prime}}^{a}\right)\tilde{O}_{l^{\prime}l}^{c}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$,
$\Omega^{a}_{l}\equiv$ $-2\varepsilon^{abc}\sum_{\gamma^{\prime}\neq\gamma}$ $\mathrm{Im}\langle\gamma|\partial^{b}_{\mathbf{k}}\hat{\mathcal{H}}|\gamma^{\prime}\rangle$ $\langle\gamma^{\prime}|\partial^{c}_{\mathbf{k}}\hat{\mathcal{H}}|\gamma\rangle$/$(\epsilon^{\gamma}_{\mathbf{k}}-\epsilon^{\gamma^{\prime}}_{\mathbf{k}})^{2}$, $\chi_{2}^{\mathrm{SSK}}=\frac{\tau^{3}e^{3}}{2\hbar}\sum_{ll^{\prime}l^{\prime\prime}}\left(v_{l^{\prime\prime}}^{a}+v_{l^{\prime}}^{a}\right)\varpi_{ll^{\prime}}^{A}\bar{O}_{l^{\prime\prime}l^{\prime}}^{c}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$,
@@
$O^{a}_{ll^{\prime}}$$\equiv$$(2\pi/\hbar)|T_{ll^{\prime}}|^{2}\delta r^{a}_{ll^{\prime}}\partial_{\epsilon_{l}}\delta(\epsilon_{l}-\epsilon_{l^{\prime}})$, $\quad\quad\quad+\frac{\tau^{3}e^{3}}{2\hbar}\sum_{ll^{\prime}}v_{l^{\prime}}^{a}v_{l}^{sj,b}\left(\partial_{\mathbf{k}}^{c}-\partial_{\mathbf{k}^{\prime}}^{c}\right)\varpi_{l^{\prime}l}^{A}\partial_{\epsilon_{l}}f_{l}^{(0)}$,
$1/\tau\approx 1/\tau_{\bf k}\equiv\sum_{\bf k^{\prime}}\varpi_{\mathbf{k}^{\prime}\mathbf{k}^{\prime}}^{(2)}\left[1-\cos\left(\phi-\phi^{\prime}\right)\right]\delta\left(\epsilon_{\mathrm{F}}-\epsilon_{l}\right)$. $\chi_{4}^{\mathrm{SSK}}=\frac{\tau^{3}e^{3}}{2\hbar^{2}}\sum_{ll^{\prime}}v_{l^{\prime}}^{sj,a}\left(\partial_{\mathbf{k}}^{c}-\partial_{\mathbf{k}^{\prime}}^{c}\right)\varpi_{l^{\prime}l}^{A}\partial_{\mathbf{k}}^{b}f_{l}^{(0)}$.
@@
Table 2: Nonzero quadratic conductivity elements $\chi_{abb}$ on the $x$-$y$ plane (defined as $J_{a}$ = $\chi_{abb}E_{b}^{2}$, $\{a,b\}\in\{x,y\}$, $abb$ is short for $\chi_{abb}$) at the order of quantum metric (all formulas in Tab. 1), can be found for only 9 out of the 122 magnetic point groups. The quantum metric dipole (QMD) Gao et al. (2014); Wang et al. (2021); Liu et al. (2021) as well as BSJ and BSK here have the anti-symmetric constraint $\chi_{abc}$ = $-\chi_{cba}$ Du et al. (2021b), so they have less nonzero elements Zhang et al. (2023); Zhu et al. (2024) than 2SJ, 2SK, and SSK found in this work. The $C_{2}$ axis is assumed in order as the $y$- and $x$- axis, the mirror plane $\sigma_{v}$ is assumed in order as the $zx$- and $yz$-plane, and the mirror plane $\sigma_{h}$ is assumed as the $x$-$y$ plane. We use the MSGCorep codes Liu et al. (2023b) to filter the groups without time-reversal and inversion but with the combined time-reversal and inversion symmetry and find their matrices.
@@
is large ($n_{i}$ is small, i.e., the clean limit), where $n_{i}$ is the impurity density, $V_{0}^{2}$ ($V_{1}^{3}$) measures the correlations of two (three) scattering events. $\chi^{\mathrm{QMD}}$, $\chi^{\mathrm{BSJ}}$, $\chi^{\mathrm{BSK}}_{2}$, and $\chi^{\mathrm{2SJ}}_{1,2,3}$ do not depend on $\lambda$, $\chi^{\mathrm{BSK}}_{1}$ is linear in $\lambda$, $\chi^{\mathrm{SSK}}_{1,2,3,4}$ has terms $\propto\lambda^{0}$ and $\lambda^{1}$, $\chi^{\mathrm{2SK}}_{1,2}$ has terms $\propto\lambda^{0}$, $\lambda^{1}$, and $\lambda^{2}$ (Tab. S5 of Sup ). The Drude contribution Sodemann and Fu (2015); Wang et al. (2021) depends on $1/(n_{i}V_{0}^{2})^{2}$. Also, for $\chi_{yxy}$ of a specific model, the disorder-induced but disorder-strength-independent contributions could overwhelms the quantum metric part Atencia et al. (2023).